# CAE-T: A Channelwise AutoEncoder with Transformer for EEG Abnormality Detection thanks: Citation: Under Review

### v1

@@ Abstract
Electroencephalogram (EEG) signals are critical for detecting abnormal brain activity, but their high dimensionality and complexity pose significant challenges for effective analysis. In this paper, we propose CAE-T, a novel framework that combines a channelwise CNN-based autoencoder with a single-head transformer classifier for efficient EEG abnormality detection. The channelwise autoencoder compresses raw EEG signals while preserving channel independence, reducing computational costs and retaining biologically meaningful features. The compressed representations are then fed into the transformer-based classifier, which efficiently models long-term dependencies to distinguish between normal and abnormal signals. Evaluated on the TUH Abnormal EEG Corpus, the proposed model achieves 85.0% accuracy, 76.2% sensitivity, and 91.2% specificity at the per-case level, outperforming baseline models such as EEGNet, Deep4Conv, and FusionCNN. Furthermore, CAE-T requires only 202M FLOPs and 2.9M parameters, making it significantly more efficient than transformer-based alternatives. The framework retains interpretability through its channelwise design, demonstrating great potential for future applications in neuroscience research and clinical practice. The source code is available at https://github.com/YossiZhao/CAE-T.
@@ 2.1 Task formulation
where $y$ represents the predicted class label (e.g., normal or abnormal EEG). Together, these stages form the backbone of the proposed CAE-T model.
2.2 Architecture of CAE-T
The CAE-T model consists of two main components: a channelwise autoencoder for feature extraction and compression, and a single-head transformer-based classifier for abnormality detection. The overall structure is depicted in Figure 1.
Figure 1: Overview of the proposed CAE-T framework for EEG-based abnormality detection. The framework includes three primary stages: (i) Preprocessing, where raw EEG signals are downsampled, segmented using a sliding window, and z-normalized to reduce noise and standardize inputs; (ii) Channelwise Autoencoder, which compresses the preprocessed signals into latent representations through independent channelwise feature extraction, preserving the natural structure of EEG data across multiple layers; and (iii) Single-Head Transformer Classifier, which utilizes scaled dot-product attention to identify patterns in the compressed representations and outputs a classification of the signals as normal or pathological. The output also has potential to support interpretability via connectivity analysis and visualization.
@@ 2.2.1 Channelwise Autoencoder
To efficiently reduce the dimensionality of raw EEG data, CAE-T employs a channelwise CNN as the core of its autoencoder. The design of our autoencoder is inspired by depthwise CNNs [20, 21, 22] and related architectures applied on EEG [23, 24]. While depthwise CNNs apply convolutions channel-wise to feature maps, our design treats each EEG channel independently, maintaining channelwise features without introducing inter-channel dependencies, as shown in Figure 2. This independence is crucial for preserving the natural structure of EEG data, which is essential for neuroscience research and clinical applications.
@@ 2.2.2 Single-Head Transformer Classifier
2.3 Integration of CAE-T for EEG Analysis
The CAE-T framework integrates the channelwise autoencoder and the single-head transformer classifier into a unified model for EEG-based abnormality detection. The autoencoder first compresses the raw EEG signal into a compact latent representation, minimizing redundant information. This representation is then input into the transformer classifier, which leverages self-attention to identify patterns critical for distinguishing between normal and abnormal signals.
By combining these components, CAE-T achieves competitive performance on benchmark datasets while maintaining a lightweight design. The model’s channelwise feature extraction and efficient transformer architecture make it particularly well-suited for applications requiring both computational efficiency and interpretability.
@@ 3 Experimental settings
This section outlines the experimental design used to evaluate the proposed CAE-T model for EEG abnormality detection, including dataset details, preprocessing steps, implementation specifics, and the baseline models used for comparison.
@@ 5.1 Trade-offs between performance and efficiency
The proposed CAE-T framework achieves an optimal trade-off between computational efficiency and classification performance, addressing the inherent challenges of long-term EEG signal processing. Compared to lightweight models such as EEGNet, which are computationally inexpensive but exhibit biased predictions, our model introduces additional computational cost justified by substantial performance improvements. Specifically, CAE-T achieves an accuracy of 85.0% at the per-case level while maintaining competitive sensitivity and specificity.
@@
The single-head transformer classifier further enhances the framework’s efficiency. While conventional multi-head attention mechanisms often lead to computationally heavy models, our single-head design retains the ability to model long-range dependencies effectively without introducing excessive parameters. Together, the autoencoder and lightweight transformer ensure that CAE-T remains practical for real-world EEG applications, including resource-constrained environments.
@@ 5.2 Interpretability of the Proposed Model
A key strength of the CAE-T framework lies in its interpretability, which is particularly valuable for clinical and neuroscience applications. The channelwise autoencoder not only reduces data dimensionality but also preserves critical spatial and temporal features, enabling insights into neural dynamics and abnormal brain activity.
@@ 5.3 Capability for Long-term EEG signal processing
Long-term EEG signals, often exceeding 2 minutes, pose unique challenges due to their size, complexity, and susceptibility to noise. Traditional deep learning approaches often struggle with such sequences, either due to computational limitations or an inability to retain relevant temporal information. Our proposed CAE-T framework effectively addresses these issues by leveraging the channelwise autoencoder to compress long sequences into compact latent representations while maintaining essential features.
The experimental results demonstrate the model’s robustness in processing extended EEG recordings. CAE-T achieves high accuracy and balanced sensitivity-specificity performance across both per-signal and per-case evaluations. These findings suggest that the framework can generalize effectively to long-term EEG signals, a critical requirement for clinical applications such as continuous brain monitoring and diagnosis of neurological disorders.
Moreover, the ability of model to handle long sequences without loss of interpretability expands its applicability to real-time EEG analysis in clinical settings. For instance, continuous EEG monitoring for seizure detection or neurological assessment often generates large volumes of data that require efficient processing. CAE-T’s lightweight design and interpretability make it a strong candidate for deployment in such scenarios, offering both computational efficiency and clinically relevant outputs.
@@ 5.4 Future Directions
The results of this study highlight the potential of CAE-T to transform EEG-based abnormality detection and neural connectivity analysis. By combining computational efficiency, interpretability, and competitive performance, the framework provides a robust solution to challenges in long-term EEG signal analysis.
@@ 6 Conclusion
This study introduced CAE-T, a novel framework for EEG-based abnormality detection that integrates a channelwise CNN-based autoencoder with a transformer-based classifier. By addressing key challenges in long-term EEG signal analysis, CAE-T demonstrates a balance between computational efficiency, interpretability, and high classification performance.
@@
CAE-T has demonstrated its ability to handle long-term EEG signals efficiently, achieving a substantial reduction in computational cost relative to standalone transformer models. Its lightweight architecture and interpretable outputs make it a practical and reliable tool for real-world applications, such as continuous monitoring in clinical settings or large-scale neuroscience studies.
Looking ahead, future work will focus on expanding the interpretability of CAE-T through analysis of self-attention similarity matrices, enabling deeper insights into functional brain connectivity. Additional efforts will explore the integration of CAE-T with multi-model datasets and its application in analyzing specific neurological disorders.
In summary, the proposed CAE-T framework provides a robust, interpretable, and efficient solution for EEG signal analysis. It paves the way for advancements in brain abnormality detection and opens new opportunities for neuroscience research and clinical diagnostics.

### v2

# CwA-T: A Channelwise AutoEncoder with Transformer for EEG Abnormality Detection thanks: Citation: Under Review

@@ Abstract
Electroencephalogram (EEG) signals are critical for detecting abnormal brain activity, but their high dimensionality and complexity pose significant challenges for effective analysis. In this paper, we propose CwA-T, a novel framework that combines a channelwise CNN-based autoencoder with a single-head transformer classifier for efficient EEG abnormality detection. The channelwise autoencoder compresses raw EEG signals while preserving channel independence, reducing computational costs and retaining biologically meaningful features. The compressed representations are then fed into the transformer-based classifier, which efficiently models long-term dependencies to distinguish between normal and abnormal signals. Evaluated on the TUH Abnormal EEG Corpus, the proposed model achieves 85.0% accuracy, 76.2% sensitivity, and 91.2% specificity at the per-case level, outperforming baseline models such as EEGNet, Deep4Conv, and FusionCNN. Furthermore, CwA-T requires only 202M FLOPs and 2.9M parameters, making it significantly more efficient than transformer-based alternatives. The framework retains interpretability through its channelwise design, demonstrating great potential for future applications in neuroscience research and clinical practice. The source code is available at https://github.com/YossiZhao/CAE-T.
@@ 2.1 Task formulation
where $y$ represents the predicted class label (e.g., normal or abnormal EEG). Together, these stages form the backbone of the proposed CwA-T model.
2.2 Architecture of CwA-T
The CwA-T model consists of two main components: a channelwise autoencoder for feature extraction and compression, and a single-head transformer-based classifier for abnormality detection. The overall structure is depicted in Figure 1.
Figure 1: Overview of the proposed CwA-T framework for EEG-based abnormality detection. The framework includes three primary stages: (i) Preprocessing, where raw EEG signals are downsampled, segmented using a sliding window, and z-normalized to reduce noise and standardize inputs; (ii) Channelwise Autoencoder, which compresses the preprocessed signals into latent representations through independent channelwise feature extraction, preserving the natural structure of EEG data across multiple layers; and (iii) Single-Head Transformer Classifier, which utilizes scaled dot-product attention to identify patterns in the compressed representations and outputs a classification of the signals as normal or pathological. The output also has potential to support interpretability via connectivity analysis and visualization.
@@ 2.2.1 Channelwise Autoencoder
To efficiently reduce the dimensionality of raw EEG data, CwA-T employs a channelwise CNN as the core of its autoencoder. The design of our autoencoder is inspired by depthwise CNNs [20, 21, 22] and related architectures applied on EEG [23, 24]. While depthwise CNNs apply convolutions channel-wise to feature maps, our design treats each EEG channel independently, maintaining channelwise features without introducing inter-channel dependencies, as shown in Figure 2. This independence is crucial for preserving the natural structure of EEG data, which is essential for neuroscience research and clinical applications.
@@ 2.2.2 Single-Head Transformer Classifier
2.3 Integration of CwA-T for EEG Analysis
The CwA-T framework integrates the channelwise autoencoder and the single-head transformer classifier into a unified model for EEG-based abnormality detection. The autoencoder first compresses the raw EEG signal into a compact latent representation, minimizing redundant information. This representation is then input into the transformer classifier, which leverages self-attention to identify patterns critical for distinguishing between normal and abnormal signals.
By combining these components, CwA-T achieves competitive performance on benchmark datasets while maintaining a lightweight design. The model’s channelwise feature extraction and efficient transformer architecture make it particularly well-suited for applications requiring both computational efficiency and interpretability.
@@ 3 Experimental settings
This section outlines the experimental design used to evaluate the proposed CwA-T model for EEG abnormality detection, including dataset details, preprocessing steps, implementation specifics, and the baseline models used for comparison.
@@ 5.1 Trade-offs between performance and efficiency
The proposed CwA-T framework achieves an optimal trade-off between computational efficiency and classification performance, addressing the inherent challenges of long-term EEG signal processing. Compared to lightweight models such as EEGNet, which are computationally inexpensive but exhibit biased predictions, our model introduces additional computational cost justified by substantial performance improvements. Specifically, CwA-T achieves an accuracy of 85.0% at the per-case level while maintaining competitive sensitivity and specificity.
@@
The single-head transformer classifier further enhances the framework’s efficiency. While conventional multi-head attention mechanisms often lead to computationally heavy models, our single-head design retains the ability to model long-range dependencies effectively without introducing excessive parameters. Together, the autoencoder and lightweight transformer ensure that CwA-T remains practical for real-world EEG applications, including resource-constrained environments.
@@ 5.2 Interpretability of the Proposed Model
A key strength of the CwA-T framework lies in its interpretability, which is particularly valuable for clinical and neuroscience applications. The channelwise autoencoder not only reduces data dimensionality but also preserves critical spatial and temporal features, enabling insights into neural dynamics and abnormal brain activity.
@@ 5.3 Capability for Long-term EEG signal processing
Long-term EEG signals, often exceeding 2 minutes, pose unique challenges due to their size, complexity, and susceptibility to noise. Traditional deep learning approaches often struggle with such sequences, either due to computational limitations or an inability to retain relevant temporal information. Our proposed CwA-T framework effectively addresses these issues by leveraging the channelwise autoencoder to compress long sequences into compact latent representations while maintaining essential features.
The experimental results demonstrate the model’s robustness in processing extended EEG recordings. CwA-T achieves high accuracy and balanced sensitivity-specificity performance across both per-signal and per-case evaluations. These findings suggest that the framework can generalize effectively to long-term EEG signals, a critical requirement for clinical applications such as continuous brain monitoring and diagnosis of neurological disorders.
Moreover, the ability of model to handle long sequences without loss of interpretability expands its applicability to real-time EEG analysis in clinical settings. For instance, continuous EEG monitoring for seizure detection or neurological assessment often generates large volumes of data that require efficient processing. CwA-T’s lightweight design and interpretability make it a strong candidate for deployment in such scenarios, offering both computational efficiency and clinically relevant outputs.
@@ 5.4 Future Directions
The results of this study highlight the potential of CwA-T to transform EEG-based abnormality detection and neural connectivity analysis. By combining computational efficiency, interpretability, and competitive performance, the framework provides a robust solution to challenges in long-term EEG signal analysis.
@@ 6 Conclusion
This study introduced CwA-T, a novel framework for EEG-based abnormality detection that integrates a channelwise CNN-based autoencoder with a transformer-based classifier. By addressing key challenges in long-term EEG signal analysis, CwA-T demonstrates a balance between computational efficiency, interpretability, and high classification performance.
@@
CwA-T has demonstrated its ability to handle long-term EEG signals efficiently, achieving a substantial reduction in computational cost relative to standalone transformer models. Its lightweight architecture and interpretable outputs make it a practical and reliable tool for real-world applications, such as continuous monitoring in clinical settings or large-scale neuroscience studies.
Looking ahead, future work will focus on expanding the interpretability of CwA-T through analysis of self-attention similarity matrices, enabling deeper insights into functional brain connectivity. Additional efforts will explore the integration of CwA-T with multi-model datasets and its application in analyzing specific neurological disorders.
In summary, the proposed CwA-T framework provides a robust, interpretable, and efficient solution for EEG signal analysis. It paves the way for advancements in brain abnormality detection and opens new opportunities for neuroscience research and clinical diagnostics.